if f(x)=x^2 + 1, find and simplify the following
(-2f(t))^2 + 5 =

Answer by ikleyn(52787)

.
if f(x)=x^2 + 1, find and simplify the following
(-2f(t))^2 + 5 =
~~~~~~~~~~~~~~~~~~

Then  (-2f(t))^2 + 5 = (-2*(t^2+1))^2 + 5 = 4*(t^2+1)^2 + 5 = 4t^4 + 8t^2 + 4 + 5 = 4t^4 + 8t^2 + 9.   ANSWER

Solved.